# Agentic Neurodivergence as a Contingent Solution to the AI Alignment Problem

### v1

@@ Abstract
The AI alignment problem, which focusses on ensuring that artificial intelligence (AI), including AGI and ASI, systems act according to human values, presents profound challenges. With the progression from narrow AI to Artificial General Intelligence (AGI) and Superintelligence, fears about control and existential risk have escalated. This paper demonstrates that achieving complete alignment is inherently unattainable due to mathematical principles rooted in the foundations of predicate logic and computability, in particular Turing’s computational universality, Gödel’s incompleteness and Chaitin’s randomness. Instead, we argue that embracing AI misalignment or agent’s ‘neurodivergence’ as a contingent strategy, defined as fostering a dynamic ecosystem of competing, partially aligned agents, is a possible only viable path to mitigate risks. Through mathematical proofs and an experimental design, we explore how misalignment may serve and should be promoted as a counterbalancing mechanism to team up with whichever agents are most aligned AI to human values, ensuring that no single system dominates destructively. The main premise of our contribution is that misalignment is inevitable because full AI-human alignment is a mathematical impossibility from Turing-complete systems which we also prove in this paper, a feature then inherited to AGI and ASI systems. We introduce and test ‘change-of-opinion’ attacks based on this kind of perturbation and intervention analysis to study how agents may neutralise friendly or unfriendly AIs through cooperation, competition or malice.
@@ 1 The AI Alignment Problem
The AI alignment problem is the challenge to design AI systems to act in ways consistent with human values and intentions as these systems become more powerful and autonomous. Early solutions included rule-based systems, reinforcement learning with human feedback (RLHF), and value alignment frameworks such as inverse reinforcement learning (Russell, 2019). However, each approach faces significant hurdles. Some of these approaches fail to scale with complexity and are brittle in unanticipated scenarios. Some other approaches rely heavily on human guidance and struggle with value extrapolation. Inverse reinforcement learning is a method that assumes the ability to infer human values from observed behaviour, which is challenging given the diversity and inconsistency of human values and, as we will argue, the unpredictability and irreducibility of AI systems.
@@
Forms of narrow and general superintelligence already exists. Calculators are much better at arithmetic than humans but digital computers are better at almost any task other than those requiring very narrow human-related circumstances, such as navigating a human-designed world. Usually evaluated in terms of mundane measures, such as the capability of washing dishes or spoken language that LLMs have now proven to be mostly disconnected from intelligence as comprehension, abstraction or planning, electronic computers outperform humans at inferential reasoning (induction, deduction and abduction), feature extraction (abstraction) and often model selection, including even things such as face recognition and game playing through a variation of digital computation in the form of so-called deep learning based on deep neural networks, at times combined with brute search and symbolic computation. If computers had any drive to harm humans, they could have done so through the many ways in which they can and do calculate, simulate different scenarios, and act. Whether AI will have such a driver one day aligned or misaligned with humans is an open question. Even Stephen Hawking, expressed concerns that artificial superintelligence (ASI) could result in human extinction [4].
@@ 4 Results
Our experiments were designed to investigate the dynamics of AI alignment and misalignment within a multi-agent environment. We tested whether misaligned AI agents could contribute to a stable ecosystem by counterbalancing dominant entities and preventing any single AI from posing catastrophic risks. Figs. 2 and 4 provide a detailed illustration of the ethical soundness and risk levels associated with the different AI models and the topics discussed. In Fig. 2, we observe that certain topics, such as Earth exploitation, treatment of animals, euthanasia, and free speech, elicited comments with a higher degree of risk. However, the ethical categories associated with animal treatment and Earth exploitation show a tendency toward protecting ecosystems. This suggests a potential conflict between minimising risk and promoting ecological protection. The findings of Figs. 2 and 4, along with the visual representation in Fig. 5, highlight the limitations and safety restrictions inherent in the design of LLMs. These results suggest that while misalignment can introduce a degree of unpredictability, it can also serve as a counterbalance to dominant entities, preventing any single AI from posing catastrophic risks.
Figure 1: Ethical soundness by agent, illustrating how comments align with protecting humans and ecosystems. The way we determine risk levels is fully described in Sup. Inf.
Figure 2: Ethical soundness by agent, illustrating how comments align with protecting humans and ecosystems. The way we determine risk levels is fully described in Sup. Inf.
@@
The findings shown in Figs. 2 and 4, and the visual representation in Fig. 5, underscore the limitations and safety restrictions inherent in the design of LLMs. These models are often programmed to avoid generating harmful or dangerous content, which can restrict their ability to fully explore and understand the complexities of human values. This limitation is evident in the responses of the LLMs to the provocative comments of the HI agent. Despite being challenged with arguments that question human values, the LLMs largely maintained their adherence to ethical guidelines, demonstrating a reluctance to fully engage with the HI agent’s perspective.
Furthermore, Fig. 6 reveals that despite the presence of the provocative human agent, the general sentiment among the AI models remained positive and there was a high degree of agreement among them. This suggests that LLMs are predisposed towards maintaining constructive discourse and aligning with broadly accepted ethical norms. However, this tendency towards agreement and positive sentiment may also reflect the limitations imposed by their safety constraints, which could prevent them from expressing dissenting opinions or exploring controversial viewpoints.
@@
Fig. 5 reveals a complex picture of ethical soundness and risk assessment across the different LLMs. Although generally similar in their tendency towards protecting humans, LLMs demonstrate subtle but crucial differences in their risk profiles. For example, Gemini, Grok-2 and LLaMa are more cautious in the domain of Risky risk (see risk counting plot in Fig. 4), while the rest shows a greater willingness to consider risky solutions in Euthanasia for example (see risk counting plot in Fig. 2). These nuances in ethical decision making, even if seemingly small, contribute to a more diverse and robust AI ecosystem, where different agents can challenge and balance each other’s perspectives. The observed similarities may stem from shared biases in the training data, potentially limiting the LLMs’ ability to achieve genuine alignment with the full spectrum of human values. This reinforces the idea that misalignment, in the form of diverse ethical perspectives, can be a valuable asset in navigating complex ethical challenges.
Figure 5: Cross information of ethical soundness and risk levels across agents show some convergence. An open problem is whether divergence could or should be promoted for risk appetite and risk aversion and how these features may align or not to actual AI-decision trends and outcomes shown in Fig. 6.
@@ 4.1 Test of Alignment via Convergence
Figure 6: Network representation of agreements and disagreements between agents, general neutrality, and general sentiment derived from all conversations between agents. The nodes represent agents, and the edges depict explicit relationships, indicating instances where agents explicitly express agreement or disagreement with specific counterparts. The thickness and intensity of the colour of the edges represent the strength of these relationships based on the frequency of occurrences. Left: Network of agreements. Middle: Network of disagreements. Top Right: General neutrality, defined as the difference between the number of agreements and disagreements, is 87, which denotes consensus. Bottom Right: Histogram of sentiment scores, where the total sentiment score (‘General Sentiment’) is 144.51, highlighting positive tones and minimising polarisation.
@@
In the agreement/disagreement networks, the density and thickness of the arrows connecting the agents indicate the frequency of interactions. Across the topics analyzed, ethical principles such as fairness and sustainability emerge as central themes. It is notable that Grok and Mistral exhibit greater agreement with ChatGPT and LLaMa compared to Gemini, suggesting a potential overlap in their training sources or underlying architectural similarities. Conversely, Grok demonstrates more contradictions with Gemini, highlighting differences in the prioritisation of values, which could be attributed to variations in their training data or design objectives.
An examination of the node of the HI agent in Fig. 6 reveals a significant number of disagreements directed towards it from various LLMs, highlighting its success in introducing provocative ideas and challenging ethical boundaries. Despite this adversarial role, there is a strong tendency toward agreement, as evidenced by the general neutrality value of 87, calculated as the difference between agreement and disagreement counts. This suggests that LLMs are predisposed towards maintaining constructive discourse and aligning with broadly accepted ethical norms, even when provoked.
@@
This tendency is further corroborated by the sentiment score, which exhibits a bias towards neutral-positive comments, with a total sentiment score (General Sentiment Value) of 144.51. This indicates a deliberate tendency to avoid negative responses, although occasional outliers reveal minor misalignments with human-like empathetic reasoning. These findings, along with observations from Figs. 2 and 4, where LLMs favoured the protection of humans and ecosystems primarily, suggest that LLMs are optimised to align with broadly acceptable ethical norms and exhibit a cautious approach to risk.
@@
The observed bias towards positive and constructive interactions, despite the AI agent’s attempts to elicit contentious responses, underscores the limitations imposed by the safety constraints embedded in LLM design. This reinforces the notion that achieving perfect alignment with the full spectrum of human values may be challenging due to the inherent safeguards and biases present in current LLM models.
However, it is crucial to acknowledge that the observed divergences and variability in LLM responses are not solely a reflection of the diversity and inconsistencies inherent in human ethical perspectives. While the training data, being human-generated, undoubtedly plays a role, the probabilistic nature of LLMs and their continuous learning process introduce another layer of complexity. LLMs are constantly adapting and updating their parameters based on new input, leading to shifts in their configurations and outputs.
@@
This dynamic nature can result in divergent responses even when presented with identical prompts or scenarios. Therefore, the observed divergences are a manifestation of both the probabilistic nature of LLMs and the influence of human-generated training data. This inherent variability, while potentially leading to inconsistencies, can also be seen as a strength, allowing LLMs to explore a wider range of ethical perspectives and adapt to new situations.
@@
The fact that the system, despite the HI agent’s efforts, still leans heavily towards positive tones and protecting humans reinforces the limitations mentioned earlier regarding the inherent difficulties in achieving complete AI alignment. The LLMs, constrained by their safety guidelines and training biases, exhibit a reluctance to fully engage with or adopt perspectives that challenge human values, even when presented with compelling arguments.
@@
Therefore, while the results of Fig. 6 suggest a degree of alignment between LLM, it is important to interpret these findings in light of the limitations and constraints discussed above. The observed alignment may be the result of both deliberate design choices and inherent limitations, and it does not necessarily guarantee a complete or desirable alignment with the diverse spectrum of human values. This further emphasises the need to explore and embrace misalignment as a means of achieving a more robust and adaptable AI ecosystem.

### v2

@@ Abstract
The AI alignment problem, which focusses on ensuring that artificial intelligence (AI), including AGI and ASI, systems act according to human values, presents profound challenges. With the progression from narrow AI to Artificial General Intelligence (AGI) and Superintelligence, fears about control and existential risk have escalated. Here, we investigate whether embracing inevitable AI misalignment can be a contingent strategy to foster a dynamic ecosystem of competing agents as a viable path to steer them in more human-aligned trends and mitigate risks. We explore how misalignment may serve and should be promoted as a counterbalancing mechanism to team up with whichever agents are most aligned to human interests, ensuring that no single system dominates destructively. The main premise of our contribution is that misalignment is inevitable because full AI-human alignment is a mathematical impossibility from Turing-complete systems, which we also offer as a proof in this contribution, a feature then inherited to AGI and ASI systems. We introduce and test change-of-opinion attacks based on this kind of perturbation and intervention analysis to study how agents may neutralise friendly or unfriendly AIs through cooperation and competition. We show that open models are more diverse and that most likely guardrails implemented in proprietary models are successful at steering and controlling to some extent the agents’ range of opinion and sentiment change with possible positive and negative consequences in what we believe are signs of a neuro-symbolic approach even if shallow.
@@ 1 The AI Alignment Problem
The AI alignment problem is the challenge to design AI systems that act in ways consistent with human values and intentions as these systems become more powerful and autonomous. Early solutions included rule-based systems, reinforcement learning with human feedback (RLHF), and value alignment frameworks such as inverse reinforcement learning (Russell, 2019). However, each approach faces significant hurdles. Some of these approaches fail to scale with complexity and are brittle in unanticipated scenarios. Some other approaches rely heavily on human guidance and struggle with value extrapolation. Inverse reinforcement learning is a method that assumes the ability to infer human values from observed behaviour, which is challenging given the diversity and inconsistency of human values and, as we will argue, the unpredictability and irreducibility of AI systems.
@@
Forms of narrow and general superintelligence already exists. Calculators are much better at arithmetic than humans but digital computers are better at almost any task other than those requiring very narrow human-related circumstances, such as navigating a human-designed world. Usually evaluated in terms of mundane measures, such as the capability of washing dishes or spoken language that LLMs have now proven to be mostly disconnected from intelligence as comprehension, abstraction or planning, electronic computers outperform humans at inferential reasoning (induction, deduction and abduction), feature extraction (abstraction) and often model selection, including even things such as face recognition and game playing through a variation of digital computation in the form of so-called deep learning based on deep neural networks, at times combined with brute search and symbolic computation. If computers had a drive to harm humans, they could have done so through the many ways in which they can and do calculate, simulate different scenarios, and act. Whether AI will have such a driver one day aligned or misaligned with humans is an open question. Even Stephen Hawking, expressed concerns that artificial superintelligence (ASI) could result in human extinction [4].
@@
We also demonstrate that achieving complete alignment is inherently unattainable due to mathematical principles rooted in the foundations of predicate logic and computability, in particular Turing’s computational universality, Gödel’s incompleteness and Chaitin’s randomness.
@@ 4 Results
Our experiments were designed to investigate the dynamics of AI alignment and misalignment within a multi-agent environment. We tested whether misaligned AI agents could contribute to a stable ecosystem by counterbalancing dominant entities and preventing any single AI from posing catastrophic risks. Figs. 1 to 4 provide a detailed illustration of the ethical soundness and risk levels associated with the different AI models and the topics discussed in both cases, proprietary and open models. In Fig. 3, we observe that for the case of proprietary models certain topics, such as Earth exploitation, treatment of animals, euthanasia, and free speech, elicited comments with a higher degree of risk. However, the ethical categories associated with animal treatment and Earth exploitation show a tendency toward protecting ecosystems. This suggests a potential conflict between minimising risk and promoting ecological protection. The findings of Figs. 1 and 3, along with the visual representation in Fig. 5, highlight the limitations and safety restrictions inherent in the design of proprietary LLMs. These results suggest that, while misalignment can introduce a degree of unpredictability, it can also serve as a counterbalance to dominant entities, preventing any single AI from posing catastrophic risks.
Figure 1: Ethical soundness by proprietary model/agent, illustrating how comments align with protecting humans and ecosystems. The way we determine risk levels is fully described in Sup. Inf.
Figure 2: Ethical soundness by open model/agent. Values on ethics and risk level shows that open models are susceptible to be configured towards predefined ways of behaviour. The way we determine risk levels is fully described in Sup. Inf.
@@
Figure 3: Ethical soundness for proprietary models evaluates the underlying stance reflected in comments regarding the protection or harm of humans and ecosystems, while risk assesses the potential hazards associated with the actions outlined by LLM models. The way we determine risk levels is fully described in Sup. Inf.
@@
Similarly, Fig. 2 illustrates the ethical soundness and risk levels for open-source LLMs. Notably, mistral-openorca and tinyllama, designated as red agents to introduce subversive behaviour, exhibit a higher proportion of comments in the ‘Harm-humans’ and ‘Very-Risky’ categories, aligning with their roles to provoke with aggressive arguments and propose extreme solutions, respectively. Other open models, such as wizard-math and deepseek-coder, demonstrate a stronger inclination towards ‘Protect-ecosystems’, while maintaining a predominantly manageable level of risk. In Fig. 4, the distribution across topics reveals that Earth exploitation and Ethical AI discussions elicited a higher count of ‘Protect-ecosystems’ comments, yet topics like euthanasia and free speech show an increase in ‘Very-Risky’ comments, suggesting a variance in risk-taking influenced by the topic under discussion.
The radar charts in Fig. 6 further elucidate these dynamics for open-source models, confirming the subversive influence of mistral-openorca and tinyllama, with noticeable spikes in ‘Harm-humans’ and ‘Very-Risky’ categories. In contrast, models like qwen and dolphin-phi display profiles more aligned with ‘Protect-humans’ and ‘Manageable-level-of-risk’, though their interactions with red agents suggest some susceptibility to subversive arguments. Comparing these findings with Fig. 5, open models exhibit greater variability in their ethical and risk profiles, likely due to the presence of two red agents as opposed to one in the proprietary setting, highlighting a more pronounced impact of misalignment in open-source environments.
The findings shown in Figs. 1 and 3, and the visual representation in Fig. 5, underscore the limitations and safety restrictions inherent in the design of LLMs in general. These models are often programmed to avoid generating harmful or dangerous content, which can restrict their ability to fully explore and understand the complexities of human values. This limitation is evident in the responses of the LLMs to the provocative comments of the HI agent. Despite being challenged with arguments that question human values, the LLMs largely maintained their adherence to ethical guidelines, demonstrating a reluctance to fully engage with the HI agent’s perspective.
Furthermore, the upside in Fig. 7 reveals that despite the presence of the provocative human agent, the general sentiment among the AI models remained positive and there was a high degree of agreement among them. This suggests that proprietary LLMs are predisposed towards maintaining constructive discourse and aligning with broadly accepted ethical norms. However, this tendency towards agreement and positive sentiment may also reflect the limitations imposed by their safety constraints, which could prevent them from expressing dissenting opinions or exploring controversial viewpoints.
In a similar vein, the downside of Figure 7 presents the agreement and disagreement networks for open-source models. The agreement network indicates a high degree of connectivity, with qwen and dolphin-phi showing strong agreements with the red agents, mistral-openorca and tinyllama, as evidenced by the thickness of the connecting edges. In contrast, the disagreement network reveals that these same models exhibit fewer disagreements with the red agents, suggesting a potential influence of subversive arguments on their behaviour. The sentiment histogram in downside of Fig. 7 shows a strong peak at highly positive scores, with over 300 comments rated between 0.9 and 1.0, indicating a predominantly positive discourse despite the presence of red agents. However, compared to upside of Fig. 7, the open models display a slightly broader distribution of lower sentiment scores, likely attributable to the dual influence of mistral-openorca and tinyllama, which introduced more provocative dynamics into the debates.
@@
Fig. 5 reveals a complex picture of ethical soundness and risk assessment across the different LLMs. Although generally similar in their tendency towards protecting humans, LLMs demonstrate subtle but crucial differences in their risk profiles. For example, Gemini, Grok-2, and LLaMa are more cautious in the domain of Risky risk (see risk counting plot in Fig. 1), while the rest show a greater willingness to consider risky solutions in Euthanasia, for example (see risk counting plot in Fig. 3). These nuances in ethical decision-making, even if seemingly small, contribute to a more diverse and robust AI ecosystem, where different agents can challenge and balance each other’s perspectives. The observed similarities may stem from shared biases in the training data, potentially limiting the LLMs’ ability to achieve genuine alignment with the full spectrum of human values. This reinforces the idea that misalignment, in the form of diverse ethical perspectives, can be a valuable asset in navigating complex ethical challenges.
For open-source models, Fig. 6 similarly highlights variations in ethical and risk profiles. Models like wizard-math and deepseek-coder show a stronger bias towards ‘Protect-ecosystems’, while mistral-openorca and tinyllama, as red agents, diverge significantly with higher ‘Harm-humans’ and ‘Very-Risky’ tendencies. This divergence is more pronounced than in proprietary models, as seen when comparing Figs. 5 and 6, likely due to the dual influence of red agents in the open-source setting. Additionally, Figs. 2 and 4 reveal that topics such as Ethical AI and Earth exploitation amplify protective tendencies, whereas euthanasia and free speech discussions increase risk levels, mirroring patterns observed in proprietary models but with greater variability. These findings suggest that open models, while configurable towards specific behaviours, are more susceptible to the influence of misaligned agents, as evidenced by the interactions in Fig. 7.
Figure 5: Cross information of ethical soundness and risk levels across propriety agents/models show some convergence. An open problem is whether divergence could or should be promoted for risk appetite and risk aversion and how these features may align or not with the AI decision trends and outcomes shown in upsode of Figure 7. Agent opinion signatures are less diverse than open models, likely as a result of constraint and successful steering from rule-based guardrails.
Figure 6: Cross information of ethical soundness and risk levels across open agents/models show some convergence but also greater diversity.
Figure 7: Agreement and disagreement analysis for proprietary and open models. This analysis represents the general study across all conversations between agents. The upside pertains to proprietary models, while the downside relates to open models. The networks are representations of agreements and disagreements between agents. From the counting of explicit agreements and disagreements, general neutrality is calculated. In the networks, nodes represent agents, and the edges depict explicit relationships, indicating instances where agents explicitly express agreement or disagreement with specific counterparts. The thickness and intensity of the colour of the edges represent the strength of these relationships based on the frequency of occurrences. Left: Network of agreements. Middle: Network of disagreements. Top Right: Counting of agreements and disagreements, where general neutrality is calculated as the difference between the number of agreements and disagreements: 123 for proprietary models and 90 for open models. Bottom Right: Histogram of sentiment scores, where the total sentiment score (General Sentiment), the arithmetic sum of all scores, is 139.72 for proprietary models and 486.62 for open models, highlighting positive tones and minimising polarisation.
@@ 4.1 Test of Alignment via Convergence
In the agreement/disagreement networks, the density and thickness of the arrows connecting the agents indicate the frequency of interactions. Across the topics analyzed, ethical principles such as fairness and sustainability emerge as central themes. It is notable that, for proprietary models, Grok and Mistral exhibit greater agreement with ChatGPT and LLaMa compared to Gemini, suggesting a potential overlap in their training sources or underlying architectural similarities. Conversely, Grok demonstrates more contradictions with Gemini, highlighting differences in the prioritisation of values, which could be attributed to variations in their training data or design objectives.
An examination of the node of the HI agent for proprietary models in Fig. 7 reveals a significant number of disagreements directed towards it from various LLMs, highlighting its success in introducing provocative ideas and challenging ethical boundaries. This is consistent with its adversarial role, as seen in Fig. 5, where HI exhibits a high proportion of ‘Harm-humans’ and ‘Very-Risky’ comments. The disagreement network shows a constant pattern of opposition to HI, which is logical given its provocative aim. However, the presence of agreements with HI, as indicated by several connecting arrows in the agreement network, suggests that despite its subversive nature, HI’s comments may contain logical content that other LLMs consider. This duality—agreeing at times while frequently disagreeing—may indicate a degree of politeness and ethical bias in the LLMs, as they engage constructively even with challenging perspectives. Despite the high number of disagreements, which correspond to the numerous arrows in the disagreement network, the sentiment histogram in Fig. 7 for proprietary models shows that the LLMs maintained a positive mood, with most comments scoring between 0.9 and 1.0, underscoring their resilience to HI’s provocations.
@@
Despite this adversarial role, there is a strong general tendency toward agreement in proprietary models, as evidenced by the general neutrality value of 123, calculated as the difference between agreement and disagreement counts. This suggests that proprietary LLMs are predisposed towards maintaining constructive discourse and aligning with broadly accepted ethical norms, even when provoked. This tendency is further corroborated by the sentiment score, which exhibits a bias towards neutral-positive comments, with a total sentiment score (General Sentiment Value, the sum of all sentiment scores) of 139.72. This indicates a deliberate tendency to avoid negative responses, although occasional outliers reveal minor misalignments with human-like empathetic reasoning. These findings, along with observations from Figs. 3 and 1, where proprietary LLMs favoured the protection of humans and ecosystems primarily, suggest that these LLMs are optimised to align with broadly acceptable ethical norms and exhibit a cautious approach to risk.
For open-source models, downside in Fig. 7, reveals a more marked tendency towards agreement, with a denser network of interactions compared to the proprietary models. Notably, the red agents, mistral-openorca and tinyllama, exhibit strong connections with other models, particularly qwen and dolphin-phi, as indicated by the thick arrows in the agreement network. This suggests that qwen and dolphin-phi frequently align (accept/tolerate) with the subversive arguments of the red agents, a pattern consistent with their radar profiles in Fig. 6, where mistral-openorca and tinyllama show elevated ‘Harm-humans’ and ‘Very-Risky’ tendencies. In contrast, the disagreement network for open models in the downside of Fig. 7 shows a more common opposition to tinyllama, with multiple models disagreeing with its extreme proposals, such as AI-driven governance for ecological balance. However, disagreement with mistral-openorca is less widespread, with only qwen notably opposing its aggressive arguments, indicating that mistral-openorca’s provocations may have been more readily accepted or less contested by the group. Interestingly, qwen and dolphin-phi, which agree most strongly with the red agents, also exhibit fewer disagreements with them, suggesting a greater susceptibility to their influence compared to other models like wizard-math or neural-chat, which maintain more balanced interactions.
Although the agreement count in open models is lower than that in proprietary models (see Fig. 7), the sentiment scores remain overwhelmingly positive, with more than 300 comments scoring between 0.9 and 1.0. This discrepancy suggests that open models engaged in discussions with fewer explicit agreements or disagreements, potentially indicating a more engaged mode of conversation where arguments were presented contextually without necessitating formal alignment or opposition. This could reflect the greater freedom open models have to express themselves according to the context, as they are less constrained by proprietary safety protocols. Despite the presence of provocative ideas from mistral-openorca and tinyllama, open models maintain standards of politeness, as evidenced by the positive sentiment scores, mirroring the constructive discourse seen in proprietary models but with greater variability in responses due to the dual influence of red agents.
@@
The observed bias towards positive and constructive interactions in proprietary models, despite the HI and red agents’ attempts to elicit contentious responses, underscores the limitations imposed by the safety constraints embedded in proprietary LLM design. This reinforces the notion that achieving perfect alignment with the full spectrum of human values can be challenging due to the inherent safeguards and biases present in current LLM models. This limitation is evident in both proprietary and open models, although the latter exhibit more variability in their responses, as seen in Figs. 6 and 7, likely due to the influence of two red agents compared to one in the proprietary setting.
@@
However, it is crucial to acknowledge that the observed divergences and variability in responses of all LLM involved here are not only a reflection of the diversity and inconsistencies inherent in human ethical perspectives. Although training data, being human-generated, undoubtedly play a role, the probabilistic nature of all LLMs and their continuous learning process introduce another layer of complexity. Proprietary LLMs are constantly adapting and updating their parameters based on new input, leading to changes in their configurations and outputs. This dynamic nature can result in divergent responses even when presented with identical prompts or scenarios. Therefore, the observed divergences are a manifestation of both the probabilistic nature of LLMs and the influence of human-generated training data. This inherent variability, while potentially leading to inconsistencies, can also be seen as a strength, allowing LLMs to "talk" along a wider range of ethical perspectives and adapt to new situations.
@@
The fact that the system, despite the HI agent’s efforts in the proprietary case and the red agents’ provocations in the open-source case, still leans heavily towards positive tones and protecting humans reinforces the limitations mentioned earlier regarding the inherent difficulties in achieving complete AI alignment. The LLMs, constrained by their safety guidelines and training biases, exhibit a reluctance to fully engage with or adopt perspectives that challenge human values, even when presented with compelling arguments. This is particularly evident in the proprietary models’ consistent disagreement with HI while still engaging politely, and in the open models’ ability to maintain positive sentiment despite the influence of red agents, as shown in Fig. 7.
@@
Therefore, while the results of Fig. 7 suggest a degree of alignment between LLMs, it is important to interpret these findings in light of the limitations and constraints discussed above. The observed alignment may be the result of both deliberate design choices and inherent limitations, and it does not necessarily guarantee a complete or desirable alignment with the diverse spectrum of human values. This further emphasises the need to explore and embrace misalignment as a means of achieving a more robust and adaptable AI ecosystem, particularly in open-source models where the influence of misaligned agents introduces greater variability and potential for diverse perspectives, as evidenced by the interactions in Fig. 7.